OpenStudy (anonymous):

Choose the correct x-intercept and y-intercept of the equation 3x + 2y = 6. x-intercept: (0, 2), y-intercept: (0, 3) x-intercept: (2, 0), y-intercept: (0, 3) x-intercept: (3, 0), y-intercept: (2, 0) x-intercept: (0, 3), y-intercept: (0, 2)

hartnn (hartnn):

to get x-intercept, put y=0 in 3x+2y= 6 to get y-intercept, put x=0 in 3x+2y= 6

OpenStudy (anonymous):

then divide

hartnn (hartnn):

divide what ? if u put y=0 in 3x+2y=6, what u get x as ?